1. Valles de Benasque y Eriste

CLLOT DE CHILL 1 h. 45 min. to 2 h.

Cllot de Chill is a small valley with a refuge located in the Eriste (Grist) valley. Once in Eriste (or Grist) take the narrow asphalt road that departs from the town

Aragonesas, goes uphill to the pipe of the power station, and there joins a forest trail, which is usually in rather good condition. At the end of the trail (1,505 m.) and opposite the Espingantosa waterfall, there is a concrete bridge that crosses the river. On the other side you will find a path to the Forcan: \underline{do not take it}. Before getting there, take another track on the other side, on the right slope of the mountain that ascends to the Cllot de Chill. After a while, there is a deviation to the right. The path goes up till you get out of the woods and see the valley and the refuge.

Beautiful walk highly recommended for the whole family.